Consider an importing with an import demand function given by p=120-q, which faces an export supply function of p=2q. The government decides to impose a tariff of $3 per units of imports
1. calculate the Domestic Price and consumption before and after the imposition of tariff.
2. What is the world price of the imports before and after the imposition of the tariff?
3. Does this country benefit from the imposition of tariffs? By how much?
1
Expert's answer
2018-03-22T10:21:07-0400
Import demand function given by p = 120 - q, q = 120 - p, export supply function of p = 2q, q = p/2.
The government decides to impose a tariff of $3 per units of imports
1. The Domestic Price and consumption before the tariff is:
Qd = Qs, Pd = Ps,
120 - q = 2q,
q = 40 units,
p = 2q = $80.
After the imposition of tariff:
Export supply is q = (p - 3)/2,
120 - p = (p - 3)/2,
240 - 2p = p - 3,
3p = 243,
p = $81,
q = 120 - p = 120 - 81 = 39 units.
2. The world price of the imports before the imposition of the tariff is: p = $80.
After the imposition: p = $81.
3. This country will not benefit from the imposition of tariffs, because the price will increase and the quantity will decrease by 1 unit.